# Active Advantage-Aligned Online Reinforcement Learning with Offline Data

### v2

@@ Appendix B Additional Related Work
In an effort to mitigate the sample complexity of online RL [36], offline RL utilizes fixed datasets to train policies without online interaction, however it can be prone to extrapolation errors that lead to overestimation of state-action values. Recent off-policy actor-critic methods [14, 26, 28, 66] seek to mitigate these issues by limiting policy learning to the scope of the dataset, thereby minimizing extrapolation error. Strategies for reducing extrapolation error include value-constrained approaches [28] that aim for conservative value estimates and policy-constrained techniques [43] that ensure the policy remains close to the observed behavior in the data. There are several works that leverage advantage estimation to guide policy improvement in purely offline RL, such as LAPO [5], A2PR [32], and A2PO [48]. However, they are not well-suited for online settings because they fail to consider the importance of “onlineness,” measured by the density ratio, to align with the needs of online RL exploration and exploitation. Additionally, they do not account for uncertainty in advantage estimation.
@@
While offline RL methods can outperform the dataset’s behavior policy, they rely entirely on static data [30]. When the dataset has comprehensive coverage, methods like FQI [1] or certainty-equivalence model learning [50] can efficiently find near-optimal policies. However, in practical scenarios with limited data coverage, policies tend to be suboptimal. One approach to addressing this suboptimality is to follow offline RL with online fine-tuning, however as discussed above, existing methods are prone to catastrophic forgetting and performance drops during fine-tuning [39]. In contrast, $A^{3}$ begins with online RL while incorporating offline data to enhance the policy, selectively leveraging offline data to facilitate online policy improvement.
Appendix C Experimental Details
@@
C.1 Additional experimental results.
We explored whether different mixtures of offline datasets can be exploited by $A^{3}$. In particular, for the D4RL locomotion halfcheetah, walker2d, hopper and ant environments in Fig. 4, mix A corresponds to having the offline dataset consisting of 100% of the -simple Minari dataset, mix B corresponds to 100% -simple and 5% -medium, while mix C corresponds to 100% -simple and 10% -medium. Those proportions were chosen due to the recognizable difference in the performance of RLPD under these different settings. In particular, we observed that all RLPD runs with the offline dataset consisting of 100% of the -simple dataset and no less than 30-40% of the -medium dataset achieve similar performance. Meanwhile at lower percentages such as 5% and 10%, there is a difference between runs of RLPD, which implies that there is significant impact from the offline dataset quality to the bootstrapping from offline transitions.
@@
g hopper mix A
h hopper mix B
i hopper mix C
j ant mix A
k ant mix B
l ant mix C
@@
Figure 4: $A^{3}$ vs RLPD vs SACfD on halfcheetah, walker2d, hopper and ant with different offline dataset mixtures. $A^{3}$ outperforms or performs comparably to RLPD across diverse black-box environments.
@@
For the Adroit environments (Fig. 5) pen, relocate and door, the -cloned dataset plays the role of the -simple dataset above, while the -expert plays the role of the -medium dataset above. The mixtures were generated similarly. $A^{3}$ robustly outperforms, or at least performs on par with, RLPD across diverse black-box environments.
a pen mix A
b pen mix B
c pen mix C
d relocate mix A
e relocate mix B
f relocate mix C
g door mix A
h door mix B
i door mix C
Figure 5: $A^{3}$ vs RLPD vs SACfD on pen, relocate and door tasks with different offline dataset mixtures. $A^{3}$ outperforms or performs comparably to RLPD across diverse black-box environments.
C.2 Additional ablation studies.
@@
Ablation on density term. Fig. 6 presents further ablation studies on the density term for $A^{3}$. We see the distinction in the effectiveness of the density term is more significant over harder tasks like antmaze-medium-play.
a halfcheetah-random
b halfcheetah-medium-replay
c antmaze-medium-play
@@
Figure 6: Ablation Studies: $A^{3}$ vs $A^{3}$ without density term.
@@
Ablation on purely online setting. Fig. 7 presents further ablation studies on $A^{3}$ interacting with the environment in a purely online manner, i.e., the algorithm does not utilize access to offline data. It is consistent throughout tested environments that $A^{3}$ is able to leverage offline data effectively, especially in harder tasks like antmaze-medium-play where purely online $A^{3}$ fails to learn in the same number of steps.
a halfcheetah-random
b halfcheetah-medium-replay
c antmaze-medium-play
Figure 7: Ablation Studies: $A^{3}$ vs purely online $A^{3}$ vs purely online SAC.
@@
Ablation on priority term. Fig. 8 presents further ablation studies on the priority term for $A^{3}$, where we compare it against the sampling strategy that solely uses TD-error as the priority term, and another that combines the density term with TD-error. The superior performance of $A^{3}$ over TD+Density over tested environments indicates that prioritizing using the advantage term achieves better performance compared to the canonical TD-error term.
a halfcheetah-random
b halfcheetah-medium-replay
c antmaze-medium-play
Figure 8: Ablation Studies: $A^{3}$ vs PER (TD) vs PER with Density (TD+Density).
@@
Training and evaluation environments. Fig. 9 presents snapshots of tested D4RL locomotion tasks: halfcheetah, walker2d, hopper, ant and antmaze. halfcheetah, walker2d, hopper and ant have dense rewards, while antmaze has sparse rewards, and all environments are equipped with continuous state and action spaces.
@@
In the halfcheetah environment, the 2D agent resembles a simplified cheetah model with a torso and lined legs, with the objective of forward locomotion and maintaining balance while maximizing speed. In the walker2d environment, the 2D humanoid agent has 2 legs and multiple joints, with the objective of stable bipedal walking without falling. In the ant environment, the agent is a 3D quadrupedal agent with multiple joints and degrees of freedom, with the objective of moving forward efficiently while maintaining balance. For all of these environments, rewards are given for velocity to encourage the agent to move forward efficiently while maintaining balance, and several offline datasets, per [12], with varying characteristics, as detailed below, were tested.
@@
Fig. 10 presents snapshots of tested Adroit manipulation tasks: pen, relocate, and door. These environments involve a simulated 28-DoF robotic arm interacting with objects in a 3D space and are characterized by sparse rewards and continuous state and action spaces.
@@
In the pen environment, the robotic arm is tasked with reorienting a pen to a target orientation in free space. In the relocate environment, the arm must pick up a ball and move it to a target position, requiring coordinated grasping and relocation of an object in 3D space. The door task involves the robotic arm grasping and pulling open a door, necessitating nuanced contact dynamics and manipulation under constraints. For all of these environments, rewards are sparse and typically only given upon task completion, increasing the exploration difficulty.
@@
In the antmaze environment, the aforementioned ant agent is placed in a maze environment and must navigate from a defined start point to a goal. Rewards are binary: $1$ for reaching the goal and 00 otherwise. Varying sizes of the maze were tested: umaze (U-shaped), medium and large; which are naturally also of increasing difficulty.
Figure 9: Environments: halfcheetah, walker2d, hopper, ant and antmaze respectively.
Figure 10: Environments: pen, relocate and door respectively.
C.3 Computing infrastructure and wall-time comparison.
We performed our experiments on a cluster that includes CPU nodes (approximately 280 cores) and GPU nodes (approximately 110 NVIDIA GPUs, ranging from Titan X to A6000, set up mostly in 4- and 8-GPU configurations). On the same cluster, the wall run time of $A^{3}$ is approximately 1.5 times the run time of regular RLPD and is comparable to Off2On.
C.4 Hyperparameters and architectures.
We list the hyperparameters used for $A^{3}$ in Table 2.

### v3

@@ Appendix C Additional Related Work
In an effort to mitigate the sample complexity of online RL (Liu et al. 2023c), offline RL utilizes fixed datasets to train policies without online interaction, however it can be prone to extrapolation errors that lead to overestimation of state-action values. Recent off-policy actor-critic methods (Fujimoto, Meger, and Precup 2019; Kostrikov, Nair, and Levine 2021; Kumar et al. 2020; Wang et al. 2020) seek to mitigate these issues by limiting policy learning to the scope of the dataset, thereby minimizing extrapolation error. Strategies for reducing extrapolation error include value-constrained approaches (Kumar et al. 2020) that aim for conservative value estimates and policy-constrained techniques (Nair et al. 2020) that ensure the policy remains close to the observed behavior in the data. There are several works that leverage advantage estimation to guide policy improvement in purely offline RL, such as LAPO (Chen et al. 2022), A2PR (Liu et al. 2024a), and A2PO (Qing et al. 2024). However, they are not well-suited for online settings because they fail to consider the importance of “onlineness,” measured by the density ratio, to align with the needs of online RL exploration and exploitation. Additionally, they do not account for uncertainty in advantage estimation.
@@
While offline RL methods can outperform the dataset’s behavior policy, they rely entirely on static data (Levine et al. 2020). When the dataset has comprehensive coverage, methods like FQI (Antos, Szepesvári, and Munos 2007) or certainty-equivalence model learning (Rathnam et al. 2023) can efficiently find near-optimal policies. However, in practical scenarios with limited data coverage, policies tend to be suboptimal. One approach to addressing this suboptimality is to follow offline RL with online fine-tuning, however as discussed above, existing methods are prone to catastrophic forgetting and performance drops during fine-tuning (Luo et al. 2023). In contrast, $A^{3}$RL begins with online RL while incorporating offline data to enhance the policy, selectively leveraging offline data to facilitate online policy improvement.
Appendix D Limitations of the prior state-of-the-art.
A drawback of RLPD, as discussed by Ball et al. (2023), lies in its symmetric random sampling method applied to both online and offline data, disregarding the significance of individual transitions for evolving quality of policy. This predefined approach to sampling can potentially lead to less than optimal policy improvements due to the omission of vital data and inefficiencies arising from the use of redundant data. Such inefficiencies fail to offer any positive contribution towards enhancing policy. To address the limitation, our research presents an innovative active data sampling technique, specifically designed to optimize the use of both online and offline data in the process of policy improvement.
Appendix E Experimental Details
@@
Additional experimental results
We explored whether different mixtures of offline datasets can be exploited by $A^{3}$RL. In particular, for the D4RL locomotion: halfcheetah, walker2d and For the Adroit: relocate environments in Fig. 5, mix A corresponds to having the offline dataset consisting of 100% of the -simple Minari dataset, mix B corresponds to 100% -simple and 5% -medium, while mix C corresponds to 100% -simple and 10% -medium. Those proportions were chosen due to the recognizable difference in the performance of RLPD under these different settings. In particular, we observed that all RLPD runs with the offline dataset consisting of 100% of the -simple dataset and no less than 30-40% of the -medium dataset achieve similar performance. Meanwhile at lower percentages such as 5% and 10%, there is a difference between runs of RLPD, which implies that there is significant impact from the offline dataset quality to the bootstrapping from offline transitions.
@@
g relocate mix A
h relocate mix B
i relocate mix C
@@
Figure 5: $A^{3}$RL vs RLPD vs SACfD on (D4RL) halfcheetah, walker2d, and (Adroit) relocate with different offline dataset mixtures. $A^{3}$RL outperforms or performs comparably to RLPD across diverse black-box environments.
@@
For the Adroit environment, relocate, the -cloned dataset plays the role of the -simple dataset above, while the -expert plays the role of the -medium dataset above. The mixtures were generated similarly. $A^{3}$RL robustly outperforms, or at least performs on par with, RLPD across diverse black-box environments.
Additional ablation studies
Ablation on density term.
@@
Fig. 6(a-c) presents further ablation studies on the density term for $A^{3}$RL. We see the distinction in the effectiveness of the density term is more significant over harder tasks like antmaze-medium-play.
a halfcheetah-random (density)
b halfcheetah-medium-replay (density)
c antmaze-medium-play (density)
d halfcheetah-random (online)
e halfcheetah-medium-replay (online)
f antmaze-medium-play (online)
g halfcheetah-random (priority)
h halfcheetah-medium-replay (priority)
i antmaze-medium-play (priority)
@@
Figure 6: Ablation Studies: (a-c): $A^{3}$RL vs $A^{3}$RL without density term. (d-f): $A^{3}$RL vs purely online $A^{3}$RL vs purely online SAC. (g-i): $A^{3}$RL vs PER (TD) vs PER with Density (TD+Density).
Ablation on purely online setting.
@@
Fig. 6(d-f) presents further ablation studies on $A^{3}$RL interacting with the environment in a purely online manner, i.e., the algorithm does not utilize access to offline data. It is consistent throughout tested environments that $A^{3}$RL is able to leverage offline data effectively, especially in harder tasks like antmaze-medium-play where purely online $A^{3}$RL fails to learn in the same number of steps.
Ablation on priority term.
@@
Fig. 6(g-i) presents further ablation studies on the priority term for $A^{3}$RL, where we compare it against the sampling strategy that solely uses TD-error as the priority term, and another that combines the density term with TD-error. The superior performance of $A^{3}$RL over TD+Density over tested environments indicates that prioritizing using the advantage term achieves better performance compared to the canonical TD-error term.
Training and evaluation environments.
@@
Fig. 7 presents snapshots of tested D4RL locomotion tasks: halfcheetah, walker2d, ant and antmaze. halfcheetah, walker2d, and ant have dense rewards, while antmaze has sparse rewards, and all environments are equipped with continuous state and action spaces.
@@
In the halfcheetah environment, the 2D agent resembles a simplified cheetah model with a torso and lined legs, with the objective of forward locomotion and maintaining balance while maximizing speed. In the walker2d environment, the 2D humanoid agent has 2 legs and multiple joints, with the objective of stable bipedal walking without falling. In the ant environment, the agent is a 3D quadrupedal agent with multiple joints and degrees of freedom, with the objective of moving forward efficiently while maintaining balance. For all of these environments, rewards are given for velocity to encourage the agent to move forward efficiently while maintaining balance, and several offline datasets, per (Fu et al. 2020), with varying characteristics, as detailed below, were tested.
@@
Fig. 7(right most) presents snapshots of tested Adroit manipulation tasks: relocate. This environment involve a simulated 28-DoF robotic arm interacting with objects in a 3D space and are characterized by sparse rewards and continuous state and action spaces.
@@
In the relocate environment, the arm must pick up a ball and move it to a target position, requiring coordinated grasping and relocation of an object in 3D space. For all of these environments, rewards are sparse and typically only given upon task completion, increasing the exploration difficulty.
@@
In the antmaze environment, the aforementioned ant agent is placed in a maze environment and must navigate from a defined start point to a goal. Rewards are binary: $1$ for reaching the goal and 0 otherwise. Varying sizes of the maze were tested: umaze (U-shaped), medium and large; which are naturally also of increasing difficulty.
Figure 7: Environments: halfcheetah, walker2d, ant, antmaze respectively and relocate.
Computing infrastructure and wall-time comparison.
We performed our experiments on a cluster that includes CPU nodes (approximately 280 cores) and GPU nodes (approximately 110 NVIDIA GPUs, ranging from Titan X to A6000, set up mostly in 4- and 8-GPU configurations). On the same cluster, the wall run time of $A^{3}$RL is approximately 1.5 times the run time of regular RLPD and is comparable to Off2On.
Hyperparameters and architectures.
We list the hyperparameters used for $A^{3}$RL in Table 2.